### v1

# Experimental investigation of turbulence and turbulent thermal diffusion in strongly inhomogeneous and anisotropic forced convection

Abstract
We investigate properties of turbulence and turbulent transport of non-inertial particles described in terms of turbulent thermal diffusion in strongly inhomogeneous and anisotropic convection forced by two turbulence generators with oscillating membrane and a steady grid in the air flow (with the Rayleigh number about $10^{8}$). Velocity field and spatial distribution of particles are measured using Particle Image Velocimetry system. The temperature distribution is measured in many locations using a temperature probe equipped with 12 E - thermocouples. In the forced convection, inhomogeneities of the mean temperature field and the particle number density in the horizontal direction in the core flow is much stronger than in the vertical direction. The mean fluid velocity structure show transition between a single-roll pattern for isothermal turbulence to double-roll patterns with increase of the temperature difference between the bottom and upper walls of the chamber. For larger temperature differences, the mean fluid velocity structure returns to a single-roll pattern. The maximum of the mean particle number density is located in the regions with minimum of the mean temperature and vise versa in agreement with the predictions based on the phenomenon of turbulent thermal diffusion.
I Introduction
Turbulent transport of particles has been studied in a number of papers due to a large variety of applications in geophysical, astrophysical and industrial flows [1, 2, 3, 4, 5, 6, 7]. Important effects in turbulent transport of particles are formation of small-scale and large-scale clusters in particle spatial distributions. In a fully developed turbulence, characteristic sizes of small-scale clusters are much smaller than the integral turbulence scale $\ell_{0}$, while large-scale clusters have sizes larger than $\ell_{0}$. Similarly, characteristic life-time of small-scale clusters are much smaller than the turbulent time $\tau_{0}$ in the integral turbulence scale, whereas life-time of large-scale clusters are of the order of turbulent diffusion time which is much larger than $\tau_{0}$ [4, 5, 6, 7, 8, 9, 10, 11].
In an isothermal inhomogeneous turbulence, one of the important mechanism of large-scale particle clusters is turbophoresis that is combined effect of particle inertia and inhomogeneity of turbulence. This phenomenon causes formation of large-scale particle clusters in the vicinity of the minimum turbulence intensity [12, 13, 14, 15, 10, 16, 17, 18].
In density or temperature stratified turbulence, large-scale particle clusters are formed due to turbulent thermal diffusion [19, 20, 7] that occurs for non-inertial and inertial particles. The latter effect causes a non-zero effective pumping velocity in the turbulent flux particles in the direction opposite to the mean temperature gradient, so that particles are accumulated in the vicinity of the mean temperature minimum.
Below we shortly discuss the mechanism of turbulent thermal diffusion of small non-inertial particles in a developed turbulence. Equation for the particle number density $n(t,{\bm{x}})$ advected by fluid velocity field ${\bm{U}}(t,{\bm{x}})$ is given by
${\partial n\over\partial t}+{\bm{\nabla}}{\bf\cdot}({\bm{U}}\,n)=D\,\Delta n,$ (1)
where $D=k_{B}\,T/(3\pi\rho\,\nu\,d_{\rm p})$ is the coefficient of the molecular (Brownian) diffusion of particles, $d_{\rm p}$ is the particle diameter, $\nu$ is the kinematic viscosity of the fluid, $T$ and $\rho$ are the fluid temperature and density and $k_{B}$ is the Boltzmann constant. We consider a low-Mach-number fluid flow with inhomogeneous density, so that the continuity equation in an anelastic approximation is given by $\mbox{\boldmath$\nabla$}{\bf\cdot}(\rho\,{\bm{U}})=0$.
We discuss large-scale clustering of particles. To this end, we apply a mean-field approach and use Reynolds averaging, so that all quantities are decomposed into the mean and fluctuating parts, and the fluctuating parts have zero mean values. Averaging Eq. (1) over an ensemble of turbulent velocity field, we arrive at the mean-field equation for the particle number density:
${\partial\overline{n}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\Big{(}\overline{\bm{U}}\,\overline{n}+\langle{\bm{u}}\,n^{\prime}\rangle\Big{)}=D\,\Delta\overline{n},$ (2)
where $\overline{n}=\langle n\rangle$ is the mean particle number density, $n^{\prime}$ are particle number density fluctuations, $\langle{\bm{u}}\,n^{\prime}\rangle$ is the turbulent flux of particles, $\bm{u}$} are velocity fluctuations, and the angular brackets denote ensemble averaging. For simplicity, we assume that mean velocity $\overline{\bm{U}}$ vanishes.
To find the particle turbulent flux, we use an equation for particle number density fluctuations $n^{\prime}$, that is obtained by subtracting Eq. (2) from Eq. (1), which yields
${\partial n^{\prime}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)-D\Delta n^{\prime}=-({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}}),$
where ${\cal Q}={\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)$ is the nonlinear term, and $({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})$ are the source terms for particle number density fluctuations. The fluid density is decomposed into the mean fluid density $\overline{\rho}$ and fluctuations $\rho^{\prime}$, and for a low-Mach-number flow, $|\rho^{\prime}|\ll\overline{\rho}$. The anelastic approximation yields $\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}}\approx-\overline{\rho}^{\,-1}\,({\bm{u}}\cdot{\bm{\nabla}})\,\overline{\rho}={\bm{u}}\cdot{\bm{\lambda}}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. Therefore, $\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})=(\overline{n}/\overline{\rho}\,)({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}=-\overline{n}\,({\bm{u}}{\bf\cdot}{\bm{\lambda}})$. The ratio of the absolute values of the nonlinear term $|{\cal Q}|$ to the diffusion term $|D\Delta n^{\prime}|$ is the Péclet number for particles, that can be estimated as ${\rm Pe}=u_{0}\,\ell_{0}/D$.
Note that nonlinear equation (3) cannot be solved exactly for arbitrary Péclet numbers. We consider a one-way coupling, and take into account the effect of the turbulent velocity on the particle number density. On the other hand, we neglect the feedback effect of the particle number density on the turbulent fluid flow. This one-way coupling approximation is valid when the spatial density of particles $n\,m_{p}$ is much smaller than the fluid density $\rho$, where $m_{p}$ is the particle mass.
We apply for simplicity the dimensional analysis to solve Eq. (3). The dimension of the left-hand side of Eq. (3) is the rate of change of particle number density fluctuations $n^{\prime}/\tau_{n^{\prime}}$, where $\tau_{n^{\prime}}$ is the characteristic time of particle number density fluctuations. For large Reynolds and Péclet numbers, the characteristic time of particle number density fluctuations $\tau_{n^{\prime}}$ can be identified with the correlation time $\tau_{0}$ of the turbulent velocity field. This implies that in the framework of the dimensional analysis, we replace the left-hand side of Eq. (3) by $n^{\prime}/\tau_{0}$, that yields:
$n^{\prime}=-\tau_{0}\,\left[({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}+\overline{n}\,({\bm{\nabla}}{\bf\cdot}{\bm{u}})\right].$ (4)
Multiplying Eq. (4) by velocity fluctuations, $u_{i}$, and averaging over an ensemble of turbulent velocity field, we obtain the turbulent flux of particles as
$\left\langle n^{\prime}\,u_{i}\right\rangle$ $=$ $-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\nabla_{j}\overline{n}-\tau_{0}\,\overline{n}\,\left\langle u_{i}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$ (5)
$\equiv$ $V_{i}^{\rm eff}\,\overline{n}-D_{ij}^{(n)}\,\nabla_{j}\overline{n}.$
The last term, $D_{ij}^{(n)}\,\nabla_{j}\overline{n}$, in Eq. (5) describes the contribution to the flux of particles caused by turbulent diffusion, where $D_{ij}^{(n)}=\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle$ is the turbulent diffusion tensor. For an isotropic turbulence $\langle u_{i}u_{j}\rangle=\delta_{ij}\,\langle{\bm{u}}^{2}\rangle/3$, so that the turbulent diffusion tensor for large Péclet numbers is given by $D_{ij}^{(n)}=D_{\rm T}\delta_{ij}$, where $D_{\rm T}=\tau_{0}\,\left\langle{\bm{u}}^{2}\right\rangle/3$ is the turbulent diffusion coefficient and $\delta_{ij}$ is the unit Kronecker tensor (i.e., the diagonal components of the tensor $\delta_{ii}=1$, and the off-diagonal components are zero) .
Figure 1: Experimental setup with the forced convective turbulence (left panel) produced by two turbulence generator with oscillating membrane and a steady grid: (1) digital CCD camera; (2) and (3) turbulence generator with oscillating membrane; (4) steady grid; (5) laser light sheet; (6) temperature probe equipped with 12 E - thermocouples; (7) heat exchanger at the bottom heated wall of the chamber; (8) heat exchanger at the top cooled wall of the chamber. Semi-exploding view of the experimental system, showing the different components assembly (right panel).
The term ${\bm{V}}^{\rm eff}\,\overline{n}$ in Eq. (5) determines the contribution to the turbulent flux of particles caused by the effective pumping velocity: ${\bm{V}}^{\rm eff}=-\tau_{0}\,\left\langle{\bm{u}}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$. Applying the anelastic approximation, ${\bm{\nabla}}{\bf\cdot}{\bm{u}}=-(1/\overline{\rho})\,({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}$, we obtain the effective pumping velocity as $V_{i}^{\rm eff}=-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\lambda_{j}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. For an isotropic turbulence, the effective pumping velocity is given by [19, 20]
${\bm{V}}^{\rm eff}=D_{\rm T}\,{{\bm{\nabla}}\overline{\rho}\over\overline{\rho}},$ (6)
and the particle turbulent flux $\langle{\bm{u}}\,n^{\prime}\rangle$ is
$\left\langle{\bm{u}}\,n^{\prime}\right\rangle={\bm{V}}^{\rm eff}\,\overline{n}-D_{\rm T}\,{\bm{\nabla}}\overline{n}.$ (7)
Let us disscuss the physics related to the effective pumping velocity ${\bm{V}}^{\rm eff}$. To this end, we use the equation of state for a perfect gas, $P=(k_{B}/m_{\mu})\,\rho\,T$, that can be also rewritten for the mean fields as $\overline{P}=(k_{B}/m_{\mu})\,\overline{\rho}\,\overline{T}$. Here $k_{B}=R/N_{A}$ is the Boltzmann constant, $R$ is the gas constant, $N_{A}$ is the Avogadro number, $\mu=m_{\mu}N_{A}$ is the molar mass, and $\overline{P}$ and $\overline{T}$ are the mean pressure and mean temperature, respectively. We assume that $\overline{\rho}\,\overline{T}\gg\langle\rho^{\prime}\,\theta\rangle$, and express the gradient of the mean fluid density in terms of the gradients of the mean fluid pressure ${\bm{\nabla}}\overline{P}$ and mean fluid temperature ${\bm{\nabla}}\overline{T}$ as ${\bm{\nabla}}\,\ln\overline{\rho}={\bm{\nabla}}\ln\overline{P}-{\bm{\nabla}}\ln\overline{T}$. For small mean pressure gradient, ${\bm{\nabla}}\,\ln\overline{\rho}\approx-{\bm{\nabla}}\ln\overline{T}$, and the effective pumping velocity for non-inertial particles is given by [19, 20]
${\bm{V}}^{\rm eff}=-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}.$ (8)
Various rigorous methods yield the result similar to Eq. (8), see, e.g., Ref. [7]. Equations (6)–(8) are valid also when the mean velocity does not vanish. Mean-field equation (2) for particle number density reads
${\partial\overline{n}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\biggl{[}\biggl{(}\overline{\bm{U}}-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}\biggr{)}\overline{n}-D_{\rm T}{\bm{\nabla}}\overline{n}\biggr{]}=0,$ (9)
where we neglect small molecular (Brownian) diffusion coefficient $D$ for large Peclet numbers. We assume that the gradients along the axis $Y$ of mean temperature, mean velocity, mean number density and turbulent diffusion coefficient are much larger than those in other directions. Thus, the steady-state solution for Eq. (9) for a zero total particle flux at the boundaries reads
${\overline{n}(Y)\over\overline{n}_{0}}=\biggl{(}{\overline{T}(Y)\over\overline{T}_{0}}\biggr{)}^{-1}\,\exp\biggl{(}\int_{0}^{Y}{\overline{U}_{y}\over D_{\rm T}}\,{\rm d}Y^{\prime}\biggr{)},$ (10)
where $\overline{n}(Y=0)=\overline{n}_{0}$ and $\overline{T}(Y=0)=\overline{T}_{0}$. Equation (10) implies that maximum particle number density is attained near the region with minimum mean fluid temperature.
Phenomenon of turbulent thermal diffusion has been studied analytically applying various theoretical methods (e.g., dimensional analysis, various $\tau$ approaches in physical and Fourier space, path integral approach, direct interaction approximation and functional approach, see [19, 20, 21, 22, 23, 24, 25, 26]). This effect has been detected in direct numerical simulations (DNS) [27, 28], observed in geophysics [29], and discussed in planetary [30] and astrophysical turbulence applications [31, 32].
Turbulent thermal diffusion has been detected in laboratory experiments with micron-size particles in temperature stratified turbulence produced by one or two oscillating grids [33, 34, 35, 26, 36, 37], or by multi-fan turbulence generator [38]. This phenomenon was even observed for nanoparticles [39] in temperature stratified turbulence. All these experiments show formation of large-scale particle clusters in the regions with the mean temperature minimum caused by turbulent thermal diffusion.
In the present paper we study in laboratory experiments properties of turbulence and a large-scale clustering of small solid particles in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow. This paper is organized as follows. In Sec.  II we describe experimental setup and measurement techniques. In Sec.  III we analyse the obtained experimental results, and in Sec. IV we outline conclusions.
II Experimental setup
Figure 2: Distributions of the mean velocity field $|\overline{U}|$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K between the bottom and upper walls of the chamber. The velocity is measured in cm/s and coordinates $Y$ and $Z$ are measured in cm.
Figure 3: Distributions of the turbulent velocity $|u^{\rm(rms)}|=[\langle u_{y}^{2}\rangle+\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K. The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
Figure 4: Distributions of the turbulent anisotropy $\langle u_{y}^{2}\rangle/\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K. The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
In this section we discuss the experimental set-up and measurement technique. We study properties of turbulence and a large-scale clustering of small solid particles described in terms of the phenomenon of turbulent thermal diffusion in experiments with a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow.
The experiments are conducted in rectangular chamber with the sizes $L_{x}\times L_{y}\times L_{z}$, where $L_{x}=L_{z}=26$ cm, $L_{y}=53$ cm. Here the axis $Z$ is in the vertical direction and the axis $Y$ is perpendicular to the oscillating membrane plain (see Fig. 1). The membrane is driven using a LABVIEW based program designed for controlling the frequency of the membrane motion (varying from 10 Hz to 2000 Hz) and the r.m.s voltage (which affects the amplitude of the membrane motion). In the experiments we mainly use the frequency of the membrane motion about 17 Hz. The air converging galvanized steel ejector (with 30 cm maximum diameter, 5 cm minimum diameter and 10 cm long) is attached to the chamber wall, so the turbulent flow produced by the oscillating membrane passes through the steady grid into the chamber.
A polypropylene cone woofer pushes air through the ejector which increases the mass flow rate and velocity of the air. The outlet of the ejector is covered by a double layered Polymerized Lactic Acid square adjustable grid (24 $\times$ 24 cm2) with square array of 2-5 mm thick bars, creating a mesh structure with the mesh size $9\times 9$ mm2, which corresponds to the porosity constant $\approx 0.74$. For instance, at the frequency of 17 Hz of the oscillating membrane and r.m.s voltage of 2  V, and with a 9 mm mesh size grid, a measured turbulent velocity of up to 20 cm/s is reached at 20 mesh size distance from the steady grid.
[…]
Figure 5: Distributions of the horizontal integral turbulent length scale $\ell_{y}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K. The integral turbulent length scale is measured in cm, and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
Figure 6: Distributions of the vertical integral turbulent length scale $\ell_{z}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K. The integral turbulent length scale is measured in cm, and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
Figure 7: Dependencies of the turbulent velocity $u_{y}^{\rm rms}(Y)$ (left panel) and the horizontal integral turbulent length scale $\ell_{y}(Y)$ (right panel) on the horizontal coordinate $Y$ in the core flow averaged over $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber. Fitting curves are shown by the solid (left part) and dashed (right part) lines. The velocity is measured in cm/s and the coordinates $Y$ is normalized by $L_{z}=26$ cm.
[…]
Figure 9: Dependence of the Reynolds number ${\rm Re}(Y)$ on the horizontal coordinate $Y$ in the core flow averaged over $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber. The coordinates $Y$ is normalized by $L_{z}=26$ cm.
Figure 10: Distributions of the ratio of the horizontal integral turbulent length scale $|\ell_{y}^{\rm fluct}|$ obtained by the Reynolds averaging to that $|\ell_{y}^{\rm stoch}|$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K. The coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
[…]
The particle spatial distribution is obtained by the PIV system using the effect of the Mie light scattering by particles [43]. In particular, the mean intensity of scattered light is determined in $43\times 64$ interrogation windows with the size $32\times 32$ pixels. This allows us to find the vertical and horizontal distributions of the intensity of the scattered light. We take into account that the light radiation energy flux scattered by small particles is given by $E_{s}\propto E_{0}\Psi(\pi d_{\rm p}/\lambda;a_{0};n)$. Here $\Psi$ is the scattering function, $d_{\rm p}$ is the particle diameter, $\lambda$ is the wavelength, $a_{0}$ is the index of refraction. The energy flux incident at the particle is $E_{0}\propto\pi d_{\rm p}^{2}/4$. When $\lambda>\pi d_{\rm p}$, the scattering function $\Psi$ is obtained from the Rayleigh’s law, $\Psi\propto d_{\rm p}^{4}$. For small $\lambda$, the scattering function $\Psi$ is independent of the particle diameter and the wavelength. In a general case, the scattering function $\Psi$ is obtained using the Mie equations [44].
In addition, we take into account that the light radiation energy flux scattered by small particles is $E_{s}\propto E_{0}\,n\,(\pi d_{\rm p}^{2}/4)$. Therefore, the scattered light energy flux incident on the charge-coupled device (CCD) camera probe is proportional to the particle number density $n$. The ratio of the scattered radiation fluxes at two locations in the flow and at the image measured with the CCD camera is equal to the ratio of the particle number densities at these two locations. The scattered light intensity $E^{T}$ obtained in a temperature-stratified turbulence is normalized by the scattered light intensity E measured in the isothermal case obtained under the same conditions. Our measurements show that using different concentrations of the incense smoke, the distribution of the scattered light intensity averaged over a vertical coordinate is independent of the particle number density in the isothermal flow. Therefore, using this normalization, we can characterize the spatial distribution of particle number density $n\propto E^{T}/E_{0}$ in the temperature stratified turbulence.
The measurement technique and data processing procedure described in this section are similar to those used by us in various experiments with turbulent convection [45, 47, 46, 39, 37] and stably stratified turbulence [48, 49, 36]. The similar measurement technique and data processing procedure in the experiments have been applied previously by us to investigate the phenomenon of turbulent thermal diffusion in a homogeneous turbulence [33, 34, 35, 26], mixing of particles in inhomogeneous turbulence [50], as well as for study of small-scale particle clustering [51].
III Experimental results
In this section we analyse the obtained experimental results in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid attached to the side walls of the chamber (see Fig. 1). In this experiments, the turbulent kinetic energy is produced by buoyancy and the membrane oscillations. In Fig. 2 we show the mean velocity field in the core flow (excluding near-boundary regions in the chamber) for isothermal turbulence (see Fig. 2a), and forced convective turbulence at the temperature differences $\Delta T=40$ K (see Fig. 2b), $\Delta T=50$ K (see Fig. 2c) and $\Delta T=60$ K (see Fig. 2d) between the bottom and upper walls of the chamber. The mean velocity patterns in these experiments are very complicated containing one or two large-scale circulations separated by jets produced by two turbulence generators with oscillating membranes and a steady grids. In particular, we observe a transition between a single-roll pattern for isothermal turbulence (Fig. 2a) to double-roll pattern with increase of the temperature differences between the bottom and upper walls of the chamber (Figs. 2b,c), and at larger temperature differences a single-roll structure is observed again (Fig. 2d). Large-scale shear in such mean velocity fields contributes to the turbulence production rate. The imposed temperature differences $\Delta T$ between the bottom and upper walls of the chamber increases buoyancy in convective turbulence which strongly affects the mean velocity patterns. The turbulence generators in the side walls of the chamber produce jets in opposite directions creating strong shear flows.
[…]
Figures 38 demonstrate that convective turbulence is strongly inhomogeneous and anisotropic. The inhomogeneity of the turbulence intensity in the horizontal $Y$ direction (along the jets) is stronger than in the vertical $Z$ direction. The isothermal turbulence and convective turbulence are very different, so the buoyancy and the mean temperature gradients strongly affect the distributions of the turbulent intensity. In particular, the isothermal turbulence is more anisotropic than the convective turbulence, i.e., large-scale circulations decreases anisotropy, especially away from jets.
In Fig. 7 and Fig. 9 we show the dependencies of the turbulent velocity $u_{y}^{\rm rms}(Y)$ (the left panel of Fig. 7), the horizontal integral turbulent length scale $\ell_{y}(Y)$ (the right panel of Fig. 7) and the Reynolds number ${\rm Re}(Y)$ (Fig. 9) on the horizontal coordinate $Y$ in the core flow averaged over vertical coordinate $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets created by the two turbulence generators) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber.
In various cases of isothermal and forced convection, we observe the scalings $u_{y}^{\rm rms}(Y)\propto Y^{-1}$ and $\ell_{y}(Y)\propto Y$ in the left and right ranges of the horizontal coordinate $Y$ (see Fig. 7). For instance, the data in Fig. 7 (right panel) are fitted to linear function (solid line is for left region up to first maximum and dashed line for left from the last maximum). This yields the following profiles for $\ell_{y}(Y)$: (a) for isothermal turbulence $\ell_{y}(Y)=9.4Y-2.1$ (left region) and $\ell_{y}(Y)=-7.4Y+12.8$ (right region); (b) for $\Delta T=40$ K , $\ell_{y}(Y)=12.2Y-3.1$ (left region) and $\ell_{y}(Y)=-7.2Y+12.7$ (right region); (c) for $\Delta T=50$ K, $\ell_{y}(Y)=3.1Y+0.1$ (left region) and $\ell_{y}(Y)=-4.6Y+8.4$ (right region); (d) for $\Delta T=60$ K, $\ell_{y}(Y)=3.6Y-0.4$ (left region) and $\ell_{y}(Y)=-6.8Y+11.4$ (right region). Similarly, the following fitted profiles are obtained for the velocity $u_{y}^{\rm rms}(Y)$: (a) for isothermal turbulence $u_{y}^{\rm rms}(Y)=2.6Y^{-1}+14.1$ (left region) and $u_{y}^{\rm rms}(Y)=1.7(1.9-Y)^{-1}+14.1$ (right region); (b) for $\Delta T=40$ K, $u_{y}^{\rm rms}(Y)=1.4Y^{-1}+5.7$ (left region) and $u_{y}^{\rm rms}(Y)=2.7(2.2-Y)^{-1}+6.3$ (right region); (c) for $\Delta T=50$ K, $u_{y}^{\rm rms}(Y)=3Y^{-1}+12$ (left region) and $u_{y}^{\rm rms}(Y)=1.1(1.8-Y)^{-1}+12$ (right region); (d) for $\Delta T=60$ K, $u_{y}^{\rm rms}(Y)=4.3Y^{-1}+9$ (left region) and $u_{y}^{\rm rms}(Y)=2.2(2.7-Y)^{-1}+8.1$ (right region).
These scalings resemble qualitatively the results of the early laboratory experiments conducted with one oscillating grid in isothermal turbulence [52, 53, 54, 55, 56, 57, 58] and forced convective turbulence [36, 37], where the r.m.s. velocity behaves as $\sqrt{\langle{\bf u^{\prime}}^{2}\rangle}\propto f\,Y^{-1}$, while the horizontal integral turbulence length scale increases linearly with the distance $Y$ from a grid.
To determine the integral turbulent length scales, we integrate the two-point correlation function of velocity fluctuations over the distance $r$ between two points which varies from zero to $r_{\rm min}$, where the correlation function attains the first minimum [59]. The reason is that the measured two-point correlation function of velocity fluctuations does not tends to zero for larger $r$ but it oscillates near some non-zero constant due to a strong influence of large-scale circulations on velocity fluctuations.
For comparison, we also apply other approach to determine the integral turbulent length scales. In particular, we use the proper orthogonal decomposition [60, 61, 62], which allows us to extract from the measured velocity field the mean field, the part of velocity fluctuations which are strongly affected by the mean field (which characterises the coherent structures) and stochastic component of velocity fluctuations. The integral scale of turbulence is determined by a stochastic component of velocity fluctuations.
In Fig. 10 we show the distributions of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$ of the horizontal integral turbulent length scale $\ell_{y}^{\rm fluct}$ obtained by usual Reynolds averaging (where velocity is decomposed into the mean and fluctuations) to that $\ell_{y}^{\rm stoch}$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow for isothermal turbulence and for forced convective turbulence at the various temperature differences between the bottom and upper walls of the chamber. As follows from Fig. 10 both methods yield the similar spatial distributions of the integral turbulent length scale and compared values of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$.
We also study the phenomenon of turbulent thermal diffusion in a forced convective turbulence by measuring the spatial distributions of the mean temperature and the mean particle number density. The initial spatial distributions of small solid particles injected into the chamber, are nearly homogeneous and isotropic. In a temperature-stratified turbulence, the spatial distributions of the mean particle number density is expected to be strongly inhomogeneous due to the effective pumping velocity ${\bm{V}}^{\rm eff}$ that is proportional $\propto-D_{T}\,{\bm{\nabla}}\ln\overline{T}$, so particles should be accumulated at the vicinity of the mean temperature minimum.
In left panels of Figs. 1113, we show the distributions of the mean temperature $\overline{T}(Y,Z)$ in the core flow for forced convective turbulence at the temperature difference $\Delta T=40$ K (see Fig. 11), $\Delta T=50$ K (see Fig. 12) and $\Delta T=60$ K (see Fig. 13) between the bottom and upper walls of the chamber. We see that the inhomogeneity of the mean temperature in the horizontal $Y$ direction (along the jets) is much more stronger than in the vertical $Z$ direction.
Figure 11: Distributions of the mean temperature $\overline{T}(Y,Z)$ (left panel) and the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$ (right panel) in the core flow for forced convective turbulence at the temperature difference $\Delta T=40$ K between the bottom and upper walls of the chamber. The coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
Figure 12: Distributions of the mean temperature $\overline{T}(Y,Z)$ (left panel) and the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$ (right panel) in the core flow for forced convective turbulence at the temperature difference $\Delta T=50$ K between the bottom and upper walls of the chamber. The coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
Figure 13: Distributions of the mean temperature $\overline{T}(Y,Z)$ (left panel) and the normalized mean particle number density $\overline{n}(Y,Z)$ (right panel) in the core flow for forced convective turbulence at the temperature difference $\Delta T=60$ K between the bottom and upper walls of the chamber. The coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm, and the mean particle number density $\overline{n}(Y,Z)$ is normalised by the maximum mean particle number density $\overline{n}_{0}$.
Figure 14: Dependencies of the mean temperature $\overline{T}(Y)/\overline{T}_{0}$ and the normalized mean particle number density $\overline{n}(Y)/\overline{n}_{0}$ on the horizontal coordinate $Y$ in the core flow for three fixed vertical positions: $Z/L_{z}=0.23$ (diamonds), $Z/L_{z}=0.4$ (circles), $Z/L_{z}=0.58$ (crosses), for forced convective turbulence at the temperature differences: (a) $\Delta T=40$ K; (b) $\Delta T=50$ K and (c) $\Delta T=60$ K between the bottom and upper walls of the chamber. The coordinates $Y$ is normalized by $L_{z}=26$ cm.
In right panels of Figs. 1113, we show the distributions of the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$. As expected, the inhomogeneity of the particle number density in the horizontal direction is much more stronger than in the vertical direction, and the maximum of the mean particle number density is located in the regions with minimum of the mean temperature $\overline{T}(Y,Z)$ in accordance with the phenomenon of turbulent thermal diffusion (see Figs. 1113).
In Fig. 14 we plot the dependencies of the mean temperature $\overline{T}(Y)/\overline{T}_{0}$ and the normalized mean particle number density $\overline{n}(Y)/\overline{n}_{0}$ on the horizontal coordinate $Y$ in the core flow for three fixed vertical positions: $Z/L_{z}=0.23$ (diamonds), $Z/L_{z}=0.4$ (circles), $Z/L_{z}=0.58$ (crosses), for forced convective turbulence at different temperature differences between the bottom and upper walls of the chamber. In spite of the presence of the large-scale circulations superimposed to the convective turbulence, Fig. 14 demonstrates the tendency of the localization of the maximum of the mean particle number density in the vicinity of the minimum in the mean temperature distribution, and vise versa.
Note that the terminal fall velocity for the micron-size particles is about $10^{-2}$ cm/s, while the turbulent velocity in the experiments with the forced convective turbulence is much larger than the particle terminal fall velocity (it is about $25$ cm/s near the turbulence generator and is more than $10$ cm/s far from the turbulence generator). Also the mean fluid velocity in several times smaller than the turbulent velocity. This implies that sedimentation of particles does not play any role in a formation of inhomogeneous particle distributions near the bottom wall of the chamber. The effective pumping velocity caused by turbulent thermal diffusion is about 1155 cm/s. This implies that the turbulent and effective pumping velocities in the experiments are much larger than the terminal fall velocity for micron-size particles.
IV Conclusions
Properties of turbulence and turbulent transport of non-inertial particles have been studied in experiments in a turbulent convection forced by two turbulence generators at the side walls of the chamber with oscillating membrane and a steady grid in the air flow with the Rayleigh number about $10^{8}$. Particle Image Velocimetry system has been used for measurements of velocity fields and mean particle number density spatial distributions, while the temperature field has been measured in many locations applying a temperature probe equipped with 12 E - thermocouples.
The observed large-scale mean flow patterns are formed by an interaction between the Rayleigh–Bénard convective cells and jets produced by the turbulence generators. The mean fluid flow patterns show transition between a single-roll structure for isothermal turbulence to double-roll mean flow structures with increase of the temperature differences $\Delta T$ between the bottom and upper walls of the chamber, and at larger temperature differences a single-roll structure is observed again.
This picture is opposite to that observed in a free convection [63], where two-roll patterns for small temperature differences $\Delta T$ is replaced by a single-roll pattern with increasing $\Delta T$. Moreover, the two-roll patterns is recovered for large temperature differences $\Delta T$ in a free convection. Turbulence in the forced convection is produced by buoyancy and large-scale velocity shear of large-scale circulations and colliding jets from the side walls of the system.
These experiments show that distributions of the mean particle number density are anti-correlated with the mean temperature distributions. In particular, the maximum particle number density is located in the regions with the minimum mean fluid temperature, which is qualitatively consistent with the predictions based on the phenomenon of turbulent thermal diffusion. Deviations from this trend occurs in regions with strong vertical mean fluid motions.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@
 Abstract
-We investigate properties of turbulence and turbulent transport of non-inertial particles described in terms of turbulent thermal diffusion in strongly inhomogeneous and anisotropic convection forced by two turbulence generators with oscillating membrane and a steady grid in the air flow (with the Rayleigh number about $10^{8}$).
+We investigate properties of turbulence and turbulent transport of non-inertial particles described in terms of turbulent thermal diffusion in strongly inhomogeneous and anisotropic convection forced by two similar turbulence generators with oscillating membrane and a steady grid in the air flow (with the Rayleigh number about $10^{8}$).
 Velocity field and spatial distribution of particles are measured using Particle Image Velocimetry system.
 The temperature distribution is measured in many locations using a temperature probe equipped with 12 E - thermocouples.
-In the forced convection, inhomogeneities of the mean temperature field and the particle number density in the horizontal direction in the core flow is much stronger than in the vertical direction.
+In the forced convection, the gradients of the mean temperature field and the particle number density in the horizontal direction in the core flow are much stronger than in the vertical direction.
 The mean fluid velocity structure show transition between a single-roll pattern for isothermal turbulence to double-roll patterns with increase of the temperature difference between the bottom and upper walls of the chamber.
 For larger temperature differences, the mean fluid velocity structure returns to a single-roll pattern.
-The maximum of the mean particle number density is located in the regions with minimum of the mean temperature and vise versa in agreement with the predictions based on the phenomenon of turbulent thermal diffusion.
+In the turbulent regions with large mean temperature gradients, the dominant effect of the large-scale particle clustering is turbulent thermal diffusion, resulting in that the maximum of the mean particle number density is located in the regions with minimum of the mean temperature and vise versa.
+Deviations from this feature is observed in the regions with strong mean fluid velocities where the mean temperature gradients are small.
 I Introduction
@@
 In an isothermal inhomogeneous turbulence, one of the important mechanism of large-scale particle clusters is turbophoresis that is combined effect of particle inertia and inhomogeneity of turbulence.
-This phenomenon causes formation of large-scale particle clusters in the vicinity of the minimum turbulence intensity [12, 13, 14, 15, 10, 16, 17, 18].
+This phenomenon causes the formation of large-scale particle clusters in the vicinity of the minimum turbulence intensity [12, 13, 14, 15, 10, 16, 17, 18].
 In density or temperature stratified turbulence, large-scale particle clusters are formed due to turbulent thermal diffusion [19, 20, 7] that occurs for non-inertial and inertial particles.
 The latter effect causes a non-zero effective pumping velocity in the turbulent flux particles in the direction opposite to the mean temperature gradient, so that particles are accumulated in the vicinity of the mean temperature minimum.
-Below we shortly discuss the mechanism of turbulent thermal diffusion of small non-inertial particles in a developed turbulence.
+Below we shortly discuss the mechanism of turbulent thermal diffusion resulting in the large-scale clustering of small non-inertial particles in a developed turbulence.
 Equation for the particle number density $n(t,{\bm{x}})$ advected by fluid velocity field ${\bm{U}}(t,{\bm{x}})$ is given by
@@
 where $D=k_{B}\,T/(3\pi\rho\,\nu\,d_{\rm p})$ is the coefficient of the molecular (Brownian) diffusion of particles, $d_{\rm p}$ is the particle diameter, $\nu$ is the kinematic viscosity of the fluid, $T$ and $\rho$ are the fluid temperature and density and $k_{B}$ is the Boltzmann constant.
-We consider a low-Mach-number fluid flow with inhomogeneous density, so that the continuity equation in an anelastic approximation is given by $\mbox{\boldmath$\nabla$}{\bf\cdot}(\rho\,{\bm{U}})=0$.
-We discuss large-scale clustering of particles.
-To this end, we apply a mean-field approach and use Reynolds averaging, so that all quantities are decomposed into the mean and fluctuating parts, and the fluctuating parts have zero mean values.
+We consider a low-Mach-number fluid flow with inhomogeneous density, so that the continuity equation in an anelastic approximation is given by $\mbox{$\nabla$}{\bf\cdot}(\rho\,{\bm{U}})=0$.
+To describe the large-scale clustering of particles, we apply a mean-field approach and use Reynolds averaging, so that all quantities are decomposed into the mean and fluctuating parts, and the fluctuating parts have zero mean values.
 Averaging Eq.
@@
 where $\overline{n}=\langle n\rangle$ is the mean particle number density, $n^{\prime}$ are particle number density fluctuations, $\langle{\bm{u}}\,n^{\prime}\rangle$ is the turbulent flux of particles, $\bm{u}$} are velocity fluctuations, and the angular brackets denote ensemble averaging.
-For simplicity, we assume that mean velocity $\overline{\bm{U}}$ vanishes.
-To find the particle turbulent flux, we use an equation for particle number density fluctuations $n^{\prime}$, that is obtained by subtracting Eq. (2) from Eq. (1), which yields
-${\partial n^{\prime}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)-D\Delta n^{\prime}=-({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}}),$
-where ${\cal Q}={\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)$ is the nonlinear term, and $({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})$ are the source terms for particle number density fluctuations.
-The fluid density is decomposed into the mean fluid density $\overline{\rho}$ and fluctuations $\rho^{\prime}$, and for a low-Mach-number flow, $|\rho^{\prime}|\ll\overline{\rho}$.
-The anelastic approximation yields $\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}}\approx-\overline{\rho}^{\,-1}\,({\bm{u}}\cdot{\bm{\nabla}})\,\overline{\rho}={\bm{u}}\cdot{\bm{\lambda}}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$.
-Therefore, $\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})=(\overline{n}/\overline{\rho}\,)({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}=-\overline{n}\,({\bm{u}}{\bf\cdot}{\bm{\lambda}})$.
-The ratio of the absolute values of the nonlinear term $|{\cal Q}|$ to the diffusion term $|D\Delta n^{\prime}|$ is the Péclet number for particles, that can be estimated as ${\rm Pe}=u_{0}\,\ell_{0}/D$.
-Note that nonlinear equation (3) cannot be solved exactly for arbitrary Péclet numbers.
-We consider a one-way coupling, and take into account the effect of the turbulent velocity on the particle number density.
-On the other hand, we neglect the feedback effect of the particle number density on the turbulent fluid flow.
-This one-way coupling approximation is valid when the spatial density of particles $n\,m_{p}$ is much smaller than the fluid density $\rho$, where $m_{p}$ is the particle mass.
-We apply for simplicity the dimensional analysis to solve Eq.
-(3).
-The dimension of the left-hand side of Eq.
-(3) is the rate of change of particle number density fluctuations $n^{\prime}/\tau_{n^{\prime}}$, where $\tau_{n^{\prime}}$ is the characteristic time of particle number density fluctuations.
-For large Reynolds and Péclet numbers, the characteristic time of particle number density fluctuations $\tau_{n^{\prime}}$ can be identified with the correlation time $\tau_{0}$ of the turbulent velocity field.
-This implies that in the framework of the dimensional analysis, we replace the left-hand side of Eq.
-(3) by $n^{\prime}/\tau_{0}$, that yields:
-$n^{\prime}=-\tau_{0}\,\left[({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}+\overline{n}\,({\bm{\nabla}}{\bf\cdot}{\bm{u}})\right].$ (4)
-Multiplying Eq. (4) by velocity fluctuations, $u_{i}$, and averaging over an ensemble of turbulent velocity field, we obtain the turbulent flux of particles as
-$\left\langle n^{\prime}\,u_{i}\right\rangle$ $=$ $-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\nabla_{j}\overline{n}-\tau_{0}\,\overline{n}\,\left\langle u_{i}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$ (5)
-$\equiv$ $V_{i}^{\rm eff}\,\overline{n}-D_{ij}^{(n)}\,\nabla_{j}\overline{n}.$
-The last term, $D_{ij}^{(n)}\,\nabla_{j}\overline{n}$, in Eq.
-(5) describes the contribution to the flux of particles caused by turbulent diffusion, where $D_{ij}^{(n)}=\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle$ is the turbulent diffusion tensor.
-For an isotropic turbulence $\langle u_{i}u_{j}\rangle=\delta_{ij}\,\langle{\bm{u}}^{2}\rangle/3$, so that the turbulent diffusion tensor for large Péclet numbers is given by $D_{ij}^{(n)}=D_{\rm T}\delta_{ij}$, where $D_{\rm T}=\tau_{0}\,\left\langle{\bm{u}}^{2}\right\rangle/3$ is the turbulent diffusion coefficient and $\delta_{ij}$ is the unit Kronecker tensor (i.e., the diagonal components of the tensor $\delta_{ii}=1$, and the off-diagonal components are zero) .
-Figure 1: Experimental setup with the forced convective turbulence (left panel) produced by two turbulence generator with oscillating membrane and a steady grid: (1) digital CCD camera; (2) and (3) turbulence generator with oscillating membrane; (4) steady grid; (5) laser light sheet; (6) temperature probe equipped with 12 E - thermocouples; (7) heat exchanger at the bottom heated wall of the chamber; (8) heat exchanger at the top cooled wall of the chamber.
-Semi-exploding view of the experimental system, showing the different components assembly (right panel).
-The term ${\bm{V}}^{\rm eff}\,\overline{n}$ in Eq.
-(5) determines the contribution to the turbulent flux of particles caused by the effective pumping velocity: ${\bm{V}}^{\rm eff}=-\tau_{0}\,\left\langle{\bm{u}}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$.
-Applying the anelastic approximation, ${\bm{\nabla}}{\bf\cdot}{\bm{u}}=-(1/\overline{\rho})\,({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}$, we obtain the effective pumping velocity as $V_{i}^{\rm eff}=-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\lambda_{j}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$.
-For an isotropic turbulence, the effective pumping velocity is given by [19, 20]
-${\bm{V}}^{\rm eff}=D_{\rm T}\,{{\bm{\nabla}}\overline{\rho}\over\overline{\rho}},$ (6)
-and the particle turbulent flux $\langle{\bm{u}}\,n^{\prime}\rangle$ is
-$\left\langle{\bm{u}}\,n^{\prime}\right\rangle={\bm{V}}^{\rm eff}\,\overline{n}-D_{\rm T}\,{\bm{\nabla}}\overline{n}.$ (7)
-Let us disscuss the physics related to the effective pumping velocity ${\bm{V}}^{\rm eff}$.
-To this end, we use the equation of state for a perfect gas, $P=(k_{B}/m_{\mu})\,\rho\,T$, that can be also rewritten for the mean fields as $\overline{P}=(k_{B}/m_{\mu})\,\overline{\rho}\,\overline{T}$.
-Here $k_{B}=R/N_{A}$ is the Boltzmann constant, $R$ is the gas constant, $N_{A}$ is the Avogadro number, $\mu=m_{\mu}N_{A}$ is the molar mass, and $\overline{P}$ and $\overline{T}$ are the mean pressure and mean temperature, respectively.
-We assume that $\overline{\rho}\,\overline{T}\gg\langle\rho^{\prime}\,\theta\rangle$, and express the gradient of the mean fluid density in terms of the gradients of the mean fluid pressure ${\bm{\nabla}}\overline{P}$ and mean fluid temperature ${\bm{\nabla}}\overline{T}$ as ${\bm{\nabla}}\,\ln\overline{\rho}={\bm{\nabla}}\ln\overline{P}-{\bm{\nabla}}\ln\overline{T}$.
-For small mean pressure gradient, ${\bm{\nabla}}\,\ln\overline{\rho}\approx-{\bm{\nabla}}\ln\overline{T}$, and the effective pumping velocity for non-inertial particles is given by [19, 20]
-${\bm{V}}^{\rm eff}=-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}.$ (8)
-Various rigorous methods yield the result similar to Eq.
-(8), see, e.g., Ref.
-[7].
-Equations (6)–(8) are valid also when the mean velocity does not vanish.
-Mean-field equation (2) for particle number density reads
+The particle turbulent flux $\langle{\bm{u}}\,n^{\prime}\rangle$ is given by [19, 20]
+$\left\langle{\bm{u}}\,n^{\prime}\right\rangle={\bm{V}}^{\rm eff}\,\overline{n}-D_{\rm T}\,{\bm{\nabla}}\overline{n},$ (3)
+where $D_{\rm T}$ is the turbulent diffusion coefficient and ${\bm{V}}^{\rm eff}$ is the effective pumping velocity that for small non-inertial particles is given by
+${\bm{V}}^{\rm eff}=D_{\rm T}\,{{\bm{\nabla}}\overline{\rho}\over\overline{\rho}}\approx-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}.$ (4)
+Here $\overline{\rho}$ and $\overline{T}$ are the mean fluid density and temperature, respectively, and we take into account that for a small mean pressure gradient, ${\bm{\nabla}}\,\ln\overline{\rho}\approx-{\bm{\nabla}}\ln\overline{T}$.
+Note that various rigorous methods yield Eqs.
+(3)–(4), which implies that this result is robust (see, e.g., Ref.
+[7]).
+For illustration, the equation for the turbulent flux of particles $\langle{\bm{u}}\,n^{\prime}\rangle$ for an isotropic turbulence is derived in Appendix.
+Using Eqs.
+(3)–(4), we rewrite Eq.
+(2) for the mean particle number density as
 ${\partial\overline{n}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\biggl[\biggl(\overline{\bm{U}}-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}\biggr)\overline{n}-D_{\rm T}{\bm{\nabla}}\overline{n}\biggr]=0,$ (5)
 where we neglect small molecular (Brownian) diffusion coefficient $D$ for large Peclet numbers.
-We assume that the gradients along the axis $Y$ of mean temperature, mean velocity, mean number density and turbulent diffusion coefficient are much larger than those in other directions.
+Note that Eqs.
+(3)–(4) are also valid when the mean velocity does not vanish.
+We assume that the gradients along the horizontal axis $Y$ of the mean temperature, mean velocity, mean number density and turbulent diffusion coefficient are much larger than those in other directions.
+This condition corresponds to that observed in the experiments discussed in this paper.
 Thus, the steady-state solution for Eq.
@@
 Equation (6) implies that maximum particle number density is attained near the region with minimum mean fluid temperature.
-Phenomenon of turbulent thermal diffusion has been studied analytically applying various theoretical methods (e.g., dimensional analysis, various $\tau$ approaches in physical and Fourier space, path integral approach, direct interaction approximation and functional approach, see [19, 20, 21, 22, 23, 24, 25, 26]).
+Figure 1: Experimental setup with the forced convective turbulence (left panel) produced by two turbulence generator with oscillating membrane and a steady grid: (1) digital CCD camera; (2) and (3) turbulence generator with oscillating membrane; (4) steady grid; (5) laser light sheet; (6) temperature probe equipped with 12 E - thermocouples; (7) heat exchanger at the bottom heated wall of the chamber; (8) heat exchanger at the top cooled wall of the chamber.
+Semi-exploding view of the experimental system, showing the different components assembly (right panel).
+Phenomenon of turbulent thermal diffusion has been studied analytically applying various theoretical methods (e.g., the dimensional analysis, the $\tau$ approaches applied in physical and Fourier spaces, the path integral approach, the direct interaction approximation and functional approach, see [7, 19, 20, 21, 22, 23, 24, 25, 26]).
 This effect has been detected in direct numerical simulations (DNS) [27, 28], observed in geophysics [29], and discussed in planetary [30] and astrophysical turbulence applications [31, 32].
@@
 This phenomenon was even observed for nanoparticles [39] in temperature stratified turbulence.
-All these experiments show formation of large-scale particle clusters in the regions with the mean temperature minimum caused by turbulent thermal diffusion.
-In the present paper we study in laboratory experiments properties of turbulence and a large-scale clustering of small solid particles in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow.
+All these experiments show the formation of large-scale particle clusters in the regions with the mean temperature minimum caused by turbulent thermal diffusion.
+In the present paper we study in laboratory experiments properties of turbulence and a large-scale clustering of small solid particles in a convective turbulence forced by two similar turbulence generators which contain both, oscillating membrane and a steady grid in the air flow.
+This forcing is completely different from that applied to turbulent convection in the previous experiments [34, 35, 37], where the large-scale circulations in turbulent convection were destroyed by the applied forcing, so that the mean velocity fields were much smaller than velocity fluctuations.
+In the present convective experiments, the forcing produces very complicated large-scale velocity field with large-scale circulations and strong horizontal jets.
+As the result, the gradients of the mean temperature and the mean particle number density in the horizontal direction occur to be much larger than in the vertical direction.
+Thus, in the first time, the large-scale clustering of small solid particles is investigated in the horizontal direction rather than in the vertical direction.
 This paper is organized as follows.
@@
 IV we outline conclusions.
+In Appendix we derive the equation for the turbulent flux of particles for an isotropic turbulence.
 II Experimental setup
-Figure 2: Distributions of the mean velocity field $|\overline{U}|$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K between the bottom and upper walls of the chamber.
-The velocity is measured in cm/s and coordinates $Y$ and $Z$ are measured in cm.
-Figure 3: Distributions of the turbulent velocity $|u^{\rm(rms)}|=[\langle u_{y}^{2}\rangle+\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
-The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
-Figure 4: Distributions of the turbulent anisotropy $\langle u_{y}^{2}\rangle/\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
-The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
 In this section we discuss the experimental set-up and measurement technique.
-We study properties of turbulence and a large-scale clustering of small solid particles described in terms of the phenomenon of turbulent thermal diffusion in experiments with a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow.
+We study properties of turbulence and a large-scale clustering of small solid particles described in terms of the phenomenon of turbulent thermal diffusion in experiments with a convective turbulence forced by two similar turbulence generators with oscillating membrane and a steady grid in the air flow.
 The experiments are conducted in rectangular chamber with the sizes $L_{x}\times L_{y}\times L_{z}$, where $L_{x}=L_{z}=26$ cm, $L_{y}=53$ cm.
@@
 A polypropylene cone woofer pushes air through the ejector which increases the mass flow rate and velocity of the air.
-The outlet of the ejector is covered by a double layered Polymerized Lactic Acid square adjustable grid (24 $\times$ 24 cm2) with square array of 2-5 mm thick bars, creating a mesh structure with the mesh size $9\times 9$ mm2, which corresponds to the porosity constant $\approx 0.74$.
+The outlet of the ejector is covered by a double layered Polymerized Lactic Acid square adjustable grid (24 $\times$ 24 cm2) with square array of 2-5 mm thick bars, creating a mesh structure with the mesh size $9\times 9$ mm2, which corresponds to the porosity constant $\approx 0.74$ (see Fig.
+1).
 For instance, at the frequency of 17 Hz of the oscillating membrane and r.m.s voltage of 2  V, and with a 9 mm mesh size grid, a measured turbulent velocity of up to 20 cm/s is reached at 20 mesh size distance from the steady grid.
@@
 […]
+Figure 2: Distributions of the mean velocity field $|\overline{U}|$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K between the bottom and upper walls of the chamber.
+The velocity is measured in cm/s and coordinates $Y$ and $Z$ are measured in cm.
+Figure 3: Distributions of the turbulent velocity $|u^{\rm(rms)}|=[\langle u_{y}^{2}\rangle+\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
+The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
+Figure 4: Distributions of the turbulent anisotropy $\langle u_{y}^{2}\rangle/\langle u_{z}^{2}\rangle]^{1/2}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
+The velocity is measured in cm/s and coordinates $Y$ and $Z$ are normalized by $L_{z}=26$ cm.
 Figure 5: Distributions of the horizontal integral turbulent length scale $\ell_{y}$ in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
@@
 […]
-Figure 9: Dependence of the Reynolds number ${\rm Re}(Y)$ on the horizontal coordinate $Y$ in the core flow averaged over $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber.
-The coordinates $Y$ is normalized by $L_{z}=26$ cm.
 Figure 9: Distributions of the ratio of the horizontal integral turbulent length scale $|\ell_{y}^{\rm fluct}|$ obtained by the Reynolds averaging to that $|\ell_{y}^{\rm stoch}|$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow: (a) for isothermal turbulence; and for forced convective turbulence at the temperature differences: (b) $\Delta T=40$ K between the bottom and upper walls of the chamber; (c) $\Delta T=50$ K; and (d) $\Delta T=60$ K.
@@
 […]
+Figure 10: Dependence of the Reynolds number ${\rm Re}(Y)$ on the horizontal coordinate $Y$ in the core flow averaged over $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber.
+The coordinates $Y$ is normalized by $L_{z}=26$ cm.
 The particle spatial distribution is obtained by the PIV system using the effect of the Mie light scattering by particles [43].
@@
 III Experimental results
-In this section we analyse the obtained experimental results in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid attached to the side walls of the chamber (see Fig.
+In this section we analyse the obtained experimental results in a convective turbulence forced by two similar turbulence generators with oscillating membrane and a steady grid attached to the side walls of the chamber (see Fig.
 1).
-In this experiments, the turbulent kinetic energy is produced by buoyancy and the membrane oscillations.
 In Fig.
@@
 2d).
-Large-scale shear in such mean velocity fields contributes to the turbulence production rate.
-The imposed temperature differences $\Delta T$ between the bottom and upper walls of the chamber increases buoyancy in convective turbulence which strongly affects the mean velocity patterns.
-The turbulence generators in the side walls of the chamber produce jets in opposite directions creating strong shear flows.
 […]
 […]
+To determine the integral turbulent length scales, we integrate the two-point correlation function of velocity fluctuations over the distance $r$ between two points which varies from zero to $r_{\rm min}$, where the correlation function attains the first minimum [52].
+The reason is that the measured two-point correlation function of velocity fluctuations does not tends to zero for larger $r$ but it oscillates near some non-zero constant due to a strong influence of large-scale circulations on velocity fluctuations.
+For comparison, we also apply other approach to determine the integral turbulent length scales.
+In particular, we use the proper orthogonal decomposition [53, 54, 55], which allows us to extract from the measured velocity field the mean field, the part of velocity fluctuations which are strongly affected by the mean field (which characterises the coherent structures) and stochastic component of velocity fluctuations.
+The integral scale of turbulence is determined by a stochastic component of velocity fluctuations.
+In Fig.
+9 we show the distributions of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$ of the horizontal integral turbulent length scale $\ell_{y}^{\rm fluct}$ obtained by usual Reynolds averaging (where velocity is decomposed into the mean and fluctuations) to that $\ell_{y}^{\rm stoch}$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow for isothermal turbulence and for forced convective turbulence at the various temperature differences between the bottom and upper walls of the chamber.
+As follows from Fig.
+9, in many regions both methods yield the similar spatial distributions of the integral turbulent length scale and compared values of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$.
+Only there are some spatial regions, where this ratio exceeds 1 substantially.
+May be this anomaly is related to the fact that in these regions there is a very complicated mean velocity field and the proper orthogonal decomposition does not work well there.
 Figures 310 demonstrate that convective turbulence is strongly inhomogeneous and anisotropic.
@@
 7) and the Reynolds number ${\rm Re}(Y)$ (Fig.
-9) on the horizontal coordinate $Y$ in the core flow averaged over vertical coordinate $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets created by the two turbulence generators) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber.
+10) on the horizontal coordinate $Y$ in the core flow averaged over vertical coordinate $Z$ in the range $0.4\leq Z/L_{z}\leq 0.8$ (in the region of jets created by the two turbulence generators) for isothermal turbulence (black) and for forced convective turbulence at the temperature differences $\Delta T=40$ K (red), $\Delta T=50$ K (blue) and $\Delta T=60$ K (green) between the bottom and upper walls of the chamber.
 In various cases of isothermal and forced convection, we observe the scalings $u_{y}^{\rm rms}(Y)\propto Y^{-1}$ and $\ell_{y}(Y)\propto Y$ in the left and right ranges of the horizontal coordinate $Y$ (see Fig.
@@
 velocity behaves as $\sqrt{\langle{\bf u^{\prime}}^{2}\rangle}\propto f\,Y^{-1}$, while the horizontal integral turbulence length scale increases linearly with the distance $Y$ from a grid.
-To determine the integral turbulent length scales, we integrate the two-point correlation function of velocity fluctuations over the distance $r$ between two points which varies from zero to $r_{\rm min}$, where the correlation function attains the first minimum [59].
-The reason is that the measured two-point correlation function of velocity fluctuations does not tends to zero for larger $r$ but it oscillates near some non-zero constant due to a strong influence of large-scale circulations on velocity fluctuations.
-For comparison, we also apply other approach to determine the integral turbulent length scales.
-In particular, we use the proper orthogonal decomposition [60, 61, 62], which allows us to extract from the measured velocity field the mean field, the part of velocity fluctuations which are strongly affected by the mean field (which characterises the coherent structures) and stochastic component of velocity fluctuations.
-The integral scale of turbulence is determined by a stochastic component of velocity fluctuations.
-In Fig.
-10 we show the distributions of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$ of the horizontal integral turbulent length scale $\ell_{y}^{\rm fluct}$ obtained by usual Reynolds averaging (where velocity is decomposed into the mean and fluctuations) to that $\ell_{y}^{\rm stoch}$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow for isothermal turbulence and for forced convective turbulence at the various temperature differences between the bottom and upper walls of the chamber.
-As follows from Fig.
-10 both methods yield the similar spatial distributions of the integral turbulent length scale and compared values of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$.
 We also study the phenomenon of turbulent thermal diffusion in a forced convective turbulence by measuring the spatial distributions of the mean temperature and the mean particle number density.
@@
 13) between the bottom and upper walls of the chamber.
-We see that the inhomogeneity of the mean temperature in the horizontal $Y$ direction (along the jets) is much more stronger than in the vertical $Z$ direction.
+We see that the gradients of the mean temperature in the horizontal $Y$ direction (along the jets) are much more stronger than in the vertical $Z$ direction.
 Figure 11: Distributions of the mean temperature $\overline{T}(Y,Z)$ (left panel) and the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$ (right panel) in the core flow for forced convective turbulence at the temperature difference $\Delta T=40$ K between the bottom and upper walls of the chamber.
@@
 1113, we show the distributions of the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$.
-As expected, the inhomogeneity of the particle number density in the horizontal direction is much more stronger than in the vertical direction, and the maximum of the mean particle number density is located in the regions with minimum of the mean temperature $\overline{T}(Y,Z)$ in accordance with the phenomenon of turbulent thermal diffusion (see Figs.
+As expected, the gradients of the particle number density in the horizontal direction are much more stronger than in the vertical direction.
+In the regions with the strong mean temperature gradients, the maximum of the mean particle number density is located in the regions with minimum of the mean temperature $\overline{T}(Y,Z)$ in accordance with the phenomenon of turbulent thermal diffusion (see Figs.
 1113).
@@
 14 demonstrates the tendency of the localization of the maximum of the mean particle number density in the vicinity of the minimum in the mean temperature distribution, and vise versa.
+However, this conclusion is only valid in the turbulent regions with large mean temperature gradients.
+For these regions, the dominant effect of large-scale particle clustering is turbulent thermal diffusion.
+Deviations from this trend occurs in the regions with strong mean fluid motions where the mean temperature gradient is small.
 Note that the terminal fall velocity for the micron-size particles is about $10^{-2}$ cm/s, while the turbulent velocity in the experiments with the forced convective turbulence is much larger than the particle terminal fall velocity (it is about $25$ cm/s near the turbulence generator and is more than $10$ cm/s far from the turbulence generator).
-Also the mean fluid velocity in several times smaller than the turbulent velocity.
-This implies that sedimentation of particles does not play any role in a formation of inhomogeneous particle distributions near the bottom wall of the chamber.
+This implies that sedimentation of particles does not play any role in a formation of inhomogeneous particle distributions.
 The effective pumping velocity caused by turbulent thermal diffusion is about 1155 cm/s.
@@
 IV Conclusions
-Properties of turbulence and turbulent transport of non-inertial particles have been studied in experiments in a turbulent convection forced by two turbulence generators at the side walls of the chamber with oscillating membrane and a steady grid in the air flow with the Rayleigh number about $10^{8}$.
+Properties of turbulence and turbulent transport of non-inertial particles have been studied in experiments in a turbulent convection forced by two similar turbulence generators at the side walls of the chamber with oscillating membrane and a steady grid in the air flow with the Rayleigh number about $10^{8}$.
 Particle Image Velocimetry system has been used for measurements of velocity fields and mean particle number density spatial distributions, while the temperature field has been measured in many locations applying a temperature probe equipped with 12 E - thermocouples.
@@
 Turbulence in the forced convection is produced by buoyancy and large-scale velocity shear of large-scale circulations and colliding jets from the side walls of the system.
-These experiments show that distributions of the mean particle number density are anti-correlated with the mean temperature distributions.
-In particular, the maximum particle number density is located in the regions with the minimum mean fluid temperature, which is qualitatively consistent with the predictions based on the phenomenon of turbulent thermal diffusion.
-Deviations from this trend occurs in regions with strong vertical mean fluid motions.
+In the turbulent regions with large mean temperature gradient, the dominant effect of large-scale particle clustering is turbulent thermal diffusion so that the distributions of the mean particle number density are anti-correlated with the mean temperature distributions.
+This implies that the maximum particle number density is located in the regions with the minimum mean fluid temperature, which is qualitatively consistent with the predictions based on the phenomenon of turbulent thermal diffusion.
+However, deviations from this trend occurs in the regions with strong mean fluid motions where the mean temperature gradient is small.
+Note that we have performed experiments with small non-inertial particles.
+For inertial particles a large-scale clustering can be more stronger, because there is an additional positive contribution to the effective pumping velocity (see, e.g., Ref.
+[7]).
+However, it is very difficult to perform experiments with inertial particles that is a subject of a separate investigation.
+In the present study we have done only a qualitative comparison between theoretical predictions of large-scale particle clustering and experimental results.
+Since we measure a two-dimensional velocity field (in the $YZ$ plane), and since the large-scale complicated three-dimensional flow contributes to the large-scale clustering of particles in the turbulent regions with small mean temperature gradients, it is not visible to perform a quantitative comparisons between theoretical and experimental results.
+ACKNOWLEDGEMENTS
+We are thankful to the two referees for providing constructive comments that improved our paper.
 AUTHOR DECLARATIONS
@@
 The data that support the findings of this study are available from the corresponding author upon reasonable request.
+Appendix A Derivation of the turbulent flux of particles
+To find the particle turbulent flux, we use an equation for particle number density fluctuations $n^{\prime}$, that is obtained by subtracting Eq. (2) from Eq. (1), which yields
+${\partial n^{\prime}\over\partial t}+{\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)-D\Delta n^{\prime}=-({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}}),$
+where ${\cal Q}={\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)$ is the nonlinear term, and $({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})$ are the source terms for particle number density fluctuations.
+Here, for simplicity, we assumed that mean velocity $\overline{\bm{U}}$ vanishes.
+The fluid density is decomposed into the mean fluid density $\overline{\rho}$ and fluctuations $\rho^{\prime}$, and for a low-Mach-number flow, $|\rho^{\prime}|\ll\overline{\rho}$.
+The anelastic approximation yields $\mbox{$\nabla$}{\bf\cdot}{\bm{u}}\approx-\overline{\rho}^{\,-1}\,({\bm{u}}\cdot{\bm{\nabla}})\,\overline{\rho}={\bm{u}}\cdot{\bm{\lambda}}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$.
+Therefore, $\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})=(\overline{n}/\overline{\rho}\,)({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}=-\overline{n}\,({\bm{u}}{\bf\cdot}{\bm{\lambda}})$.
+The ratio of the absolute values of the nonlinear term $|{\cal Q}|$ to the diffusion term $|D\Delta n^{\prime}|$ is the Péclet number for particles, that can be estimated as ${\rm Pe}=u_{0}\,\ell_{0}/D$.
+Note that nonlinear equation (7) cannot be solved exactly for arbitrary Péclet numbers.
+We consider a one-way coupling, and take into account the effect of the turbulent velocity on the particle number density.
+On the other hand, we neglect the feedback effect of the particle number density on the turbulent fluid flow.
+This one-way coupling approximation is valid when the spatial density of particles $n\,m_{p}$ is much smaller than the fluid density $\rho$, where $m_{p}$ is the particle mass.
+We apply for simplicity the dimensional analysis to solve Eq.
+(7).
+The dimension of the left-hand side of Eq.
+(7) is the rate of change of particle number density fluctuations $n^{\prime}/\tau_{n^{\prime}}$, where $\tau_{n^{\prime}}$ is the characteristic time of particle number density fluctuations.
+For large Reynolds and Péclet numbers, the characteristic time of particle number density fluctuations $\tau_{n^{\prime}}$ can be identified with the correlation time $\tau_{0}$ of the turbulent velocity field.
+This implies that in the framework of the dimensional analysis, we replace the left-hand side of Eq.
+(7) by $n^{\prime}/\tau_{0}$, that yields:
+$n^{\prime}=-\tau_{0}\,\left[({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}+\overline{n}\,({\bm{\nabla}}{\bf\cdot}{\bm{u}})\right].$ (8)
+Multiplying Eq. (8) by velocity fluctuations, $u_{i}$, and averaging over an ensemble of turbulent velocity field, we obtain the turbulent flux of particles as
+$\left\langle n^{\prime}\,u_{i}\right\rangle$ $=$ $-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\nabla_{j}\overline{n}-\tau_{0}\,\overline{n}\,\left\langle u_{i}(\mbox{$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$ (9)
+$\equiv$ $V_{i}^{\rm eff}\,\overline{n}-D_{ij}^{(n)}\,\nabla_{j}\overline{n}.$
+The last term, $D_{ij}^{(n)}\,\nabla_{j}\overline{n}$, in Eq.
+(9) describes the contribution to the flux of particles caused by turbulent diffusion, where $D_{ij}^{(n)}=\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle$ is the turbulent diffusion tensor.
+For an isotropic turbulence $\langle u_{i}u_{j}\rangle=\delta_{ij}\,\langle{\bm{u}}^{2}\rangle/3$, so that the turbulent diffusion tensor for large Péclet numbers is given by $D_{ij}^{(n)}=D_{\rm T}\delta_{ij}$, where $D_{\rm T}=\tau_{0}\,\left\langle{\bm{u}}^{2}\right\rangle/3$ is the turbulent diffusion coefficient and $\delta_{ij}$ is the unit Kronecker tensor (i.e., the diagonal components of the tensor $\delta_{ii}=1$, and the off-diagonal components are zero).
+The term ${\bm{V}}^{\rm eff}\,\overline{n}$ in Eq.
+(9) determines the contribution to the turbulent flux of particles caused by the effective pumping velocity: ${\bm{V}}^{\rm eff}=-\tau_{0}\,\left\langle{\bm{u}}(\mbox{$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$.
+Applying the anelastic approximation, ${\bm{\nabla}}{\bf\cdot}{\bm{u}}=-(1/\overline{\rho})\,({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}$, we obtain the effective pumping velocity as $V_{i}^{\rm eff}=-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\lambda_{j}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$.
+For an isotropic turbulence, the effective pumping velocity is given by [19, 20]
+${\bm{V}}^{\rm eff}=D_{\rm T}\,{{\bm{\nabla}}\overline{\rho}\over\overline{\rho}},$ (10)
+and the particle turbulent flux $\langle{\bm{u}}\,n^{\prime}\rangle$ is
+$\left\langle{\bm{u}}\,n^{\prime}\right\rangle={\bm{V}}^{\rm eff}\,\overline{n}-D_{\rm T}\,{\bm{\nabla}}\overline{n}.$ (11)
+Let us disscuss the physics related to the effective pumping velocity ${\bm{V}}^{\rm eff}$.
+To this end, we use the equation of state for a perfect gas, $P=(k_{B}/m_{\mu})\,\rho\,T$, that can be also rewritten for the mean fields as $\overline{P}=(k_{B}/m_{\mu})\,\overline{\rho}\,\overline{T}$.
+Here $k_{B}=R/N_{A}$ is the Boltzmann constant, $R$ is the gas constant, $N_{A}$ is the Avogadro number, $\mu=m_{\mu}N_{A}$ is the molar mass, and $\overline{P}$ and $\overline{T}$ are the mean pressure and mean temperature, respectively.
+We assume that $\overline{\rho}\,\overline{T}\gg\langle\rho^{\prime}\,\theta\rangle$, and express the gradient of the mean fluid density in terms of the gradients of the mean fluid pressure ${\bm{\nabla}}\overline{P}$ and mean fluid temperature ${\bm{\nabla}}\overline{T}$ as ${\bm{\nabla}}\,\ln\overline{\rho}={\bm{\nabla}}\ln\overline{P}-{\bm{\nabla}}\ln\overline{T}$.
+For small mean pressure gradient, ${\bm{\nabla}}\,\ln\overline{\rho}\approx-{\bm{\nabla}}\ln\overline{T}$, and the effective pumping velocity for non-inertial particles is given by [19, 20]
+${\bm{V}}^{\rm eff}=-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}.$ (12)